A boy cycles from his house to a friend's house at 20 km/h and home again 25km/h. If his round trip takes 9/10 of an hour, how far is it to his friend's house?
I need the working out steps
thank you

Answer by mananth(16949)

let the distance be x km
time forward = x/20 hours
...
time return = x/25 hours
...
x/20 + x/25 = 9/10
...
(5x+4x) / 100 = 9/10
9x/100 = 9/10
90x= 900
x=10 km distance to friend's house
..
m.ananth@hotmail.ca